OpenStudy (anonymous):

Find the nonpermissible replacement for b in this expression b^2/-5b+5

OpenStudy (chaise):

-5b +5 = 0 -5b = -5 b = -1 For this equation to b cannot equal -1.

OpenStudy (chaise):

b cannot equal 1*

OpenStudy (amistre64):

if only there was some rule for determining what is nonpermissible and that could apply to ALL equations ... hmmm ;)

OpenStudy (anonymous):

so it would be 1 or -1?

OpenStudy (amistre64):

what do you get for the denominator when b=-1?

OpenStudy (chaise):

it would be 1. sorry, it's late at night. You are aimining to get the denominator equal to 0. If b =1 the denominator is 0

OpenStudy (anonymous):

10

OpenStudy (anonymous):

n yea i just got it itz 1

OpenStudy (amistre64):

very good :)